Suppose for a commodity X, demand and supply curves

are given by the equations

(1)

Q d = 4 P

(2)

Q s = 2+P

(i) What is the equilibrium price?

(ii) What is the equilibrium quantity?



1
Expert's answer
2021-10-31T18:27:07-0400

At equilibrium Qd = Qs

But,

  • Qd = 4 P
  • Qs = 2+P

Hence,

  • 4P = 2+P
  • 4P - p = 2
  • 3P = 2
  • P =2/3

Equilibrium price = $2/3 or $0.67

Substitute 2/3 into either the demand equation or supply equation to get the equilibrium quantity.

  • Qd = 4P

But,

  • P =2/3

Hence,

  • Qd = 4(2/3)
  • Qd = 2.67

Equilibrium quantity = 2.67 units